OpenStudy (anonymous):

how to solve this inequality?? please help!! and show plots! x≤-11y

OpenStudy (anonymous):

One way is to plot it. Input a value for x and then work the equation out.

OpenStudy (anonymous):

To solve for y y = -x/11 Input a value for x

OpenStudy (anonymous):

so the plots would be (,0-11) (-11,0) or am I wrong?

OpenStudy (anonymous):

Input x lets say x =1 1 = -11y = -1/11=y = y = -1/11 When ever x is 1 y will = -1/11 Just keep ploting the points. First point will be (1,-1/11)

OpenStudy (anonymous):

One you have your points plotted an you can draw a line, you can test where to fill for the inequality. Here is a wolframe link to show you what it will look like http://www.wolframalpha.com/input/?i=x ≤-11y

OpenStudy (anonymous):

okay thanks!